OpenStudy (ammarah):

When the angle of elevation of the sun is 58 degree the shadow cast by a tree is 28ft. long. How tall is the tree?

OpenStudy (anonymous):

what grade are you in

OpenStudy (ammarah):

@triciaal

OpenStudy (compassionate):

 Remember SOHCAHTOA

OpenStudy (anonymous):

If the angle of inclination is 30 degrees and the length of the shadow is 12 ft, from this relation we can see that tan30=height of tree/length of shadow Height of tree = tan30*length of shadow = 6.93ft approximat…

OpenStudy (compassionate):

\[\tan(x) = \frac{ opposite }{ adjacent }\]

OpenStudy (ammarah):

so the answer would be 44.8

OpenStudy (compassionate):

\[\tan(58) = \frac{ x }{ 28 }\] \[\tan (1.6003345290410503553267330811834) = \frac{ x }{ 28 } \times \frac{ 28 }{ x }\] \[\tan(1.6003345290410503553267330811834) \times 28 = x\]\[x = 44.8093668131, ~~~ x = 44.8\]

OpenStudy (compassionate):

You did a great job!